OpenStudy (anonymous):

Area of a Trapezoid Solve for b1: A=1/2h(b1+b2) please show the steps thank you

OpenStudy (anonymous):

multiply by 2 to clear denominators\[A=h(b _{1}+b _{2})\]you could distribute the h\[A=hb _{1}+hb _{2}\]subtract the term with b2\[A-hb _{2}=hb _{1}\]divide by h\[\frac{A-hb _{1}}{h}=b _{1}\]